Show that
Rank(ST)=Rank s ,if T is non singular
Where S,T : V->V are linear transformation of a finite dimensional vector space.
1
Expert's answer
2020-12-16T11:48:49-0500

Let "S" is represented by matrix "A(m\times n)" and "T" is represented by matrix "B(n\times n)" (since "T" is non singular). Then "S\circ T" is represented by matrix "AB".

From properties of rank: if "B" is matrix of rank "n", then

"rank(AB)=rank(A)"

Proof:

let C=AB, then

"C_{ij}=\Sigma A_{ik}B_{kj}\implies rank(C)\leq rank(A)"

"A=CB^{-1}"

if "det(B)\neq0" , then "rank(C)=rank(A)"


In our case: "B" is matrix of rank "n" (since "det(B)\neq0" )

So, we get: "rank(S\circ T)=rank(S)"